Find the equation in the form shown below for each line with the given coordinates.
(-7,4) and (8,2)
Equation in slope-intercept form =

Answer by mananth(16946)

(-7,4) and (8,2)
slope = m = y2-y1/x2-x1
m= (2-4)/(8-(-7)
m=-2/15
...
y= mx +b
plug m,&(-7,4)
..
4=-2/15*-7+b
4-14/15=b
46/15=b
...
y=(-2x/15)+46/15